# Cell bulging and extrusion in a three-dimensional bubbly vertex model for curved epithelial sheets

### v1

@@ IV.1 Extrusion of a bubbly defect cell in a mean-field vertex model tissue
After benchmarking our theoretical model by comparison with the unbulged classical vertex model, we return to the bubbly formulation of the vertex model. We now drop the assumption $h_{S}^{b}=0=h_{S}^{a}$ and consider a bulged defect cell in a vertex model mean-field tissue. In the following we will assume the mean-field tissue parameters to be from the classical (non-bubbly) VM, thus only allowing for bulging of the defect cell. In principle, however, mean-field parameters could be derived for the BVM in a similar fashion as for the VM, or could be taken from experimental measurements for the tissue properties.
@@
Figure 6: Energy landscapes in configuration space and the resulting defect cell shapes. (a-c) Energy landscapes with coordinates $(\delta/2,b)$ for different surface tensions $\Gamma$. The black line marks the vertex model limit without bulging of the defect cell and the red circle marks the minimum in this limit. The dashed line is the projection onto the $(\Delta E,\delta/2)$ plane of this limit. The color codes the energy difference $\Delta E$ with respect to the reference energy, which is the vertex model minimum. The corners of the landscape at $(\delta/2,b)=(0,0)$ and $(\delta_{\mathrm{max}}/2,0)$ correspond to the extruded and maximally bent situations, respectively, cf. Fig. 3. (d) Defect opening angles $\delta/2$ (top) and basal vs. apical area ratio $A_{\mathrm{b}}/A_{\mathrm{a}}$ (bottom) of the minimal energy configuration (solid), starting from the VM reference, and of the extruded configuration (dashed). (e) Corresponding energy differences from the VM reference $\Delta E$. (f) For the bubbly vertex model with interfacial bending rigidity the defect cell is extruded for small apico-basal tension $\Gamma=0.5$, and partially extruded for large tension $\Gamma=1.2$. Shells of Caspar-Klug size $(T_{1},T_{2})=(12,0)$ with bending rigidity $\kappa_{\mathrm{if}}=100$ in dimensionless units.
@@
Again, we consider the total energy as the sum of the contributions from the defect, the cap and the outer cone region. We now, however, consider the energy difference to a reference state to quantify the energy release due to bulging. For this we consider the energy difference $\Delta E$ from the vertex model case with minimal energy, which we obtain from the full defect model in this case, cf. Fig. 5.
@@
As before, we find that for all configurations characterized by $\delta/2$ and $b$ the energy has a unique minimum in the tissue buckling radius $R_{B}$. We assume this radius to be minimal for every defect cell shape and thus can describe all possible shapes by the variables $(\delta/2,b)$, which constitutes the configuration space. Both parameters are bounded from below by 00, where $(0,0)$ is possible and corresponds to the fully extruded case, and from above by the volume constraint of the defect cell.
IV.2 Pentagonal defect shape and extrusion for identical cell properties
@@
Fig. 6 shows energy landscapes in configuration space for different surface tensions $\Gamma$. For now we assume identical cell properties, i.e. the tension in the defect cell matches the one of the surrounding tissue. We notice that, by bulging outward at the defect, the energy of the tissue can be reduced considerably in comparison to the classical vertex model reference case. For small $\Gamma$ more energy can be released by bulging outward than for larger $\Gamma$. For $\Gamma>0.5$ we observe that two minima form in the energy, one in the extruded case $(\delta/2,b)=(0,0)$, and one in a bulged case with very small $b$. The energy minimum in extrusion is higher than the bulged one for larger $\Gamma$ and an energy barrier has to be overcome from both the reference and the global minimum configuration to extrude the defect.
@@
We now numerically determined the (local) minimal energy configurations for different $\Gamma$, starting from the VM reference state. Fig. 6(d,e) depicts the resulting minimal energy configurations. Above a critical tension $\Gamma$ the luminal (basal) interface collapses but the angle remains finite, indicating a completely bulged state. The fully extruded state at $(\delta/2,b)=(0,0)$ always represents a possible configuration, but its energy increases with $\Gamma$. Looking at the energy difference from the VM state, a configuration with interfacial curvature is always energetically more favorable, cf. Fig. 6(e). This is not surprising as the surface area can be minimized by the cell becoming more spherical. The extruded case, on the other hand, requires more energy for large apico-basal tension, as we need to wrap the cell in the energetically (relatively) expensive apical interface. Lowering the tension allows for an energy release in the extruded case, which can become lower than in the bulged state. Note, however, that a bulged state persists beyond the point at which the extruded case is energetically more favorable and that an energy barrier exists in configuration space, which prevents the cell from being directly extruded from the reference VM configuration, cf. Fig. 6(b).
@@
Considering icosahedral shells in the BVM and adding a bending rigidity across triangulation points of hexagonal cells’ interfaces enables a numerical simulation similar to our continuum approach of a bubbly defect cell surrounded by VM cells. For this we consider an additional interfacial bending energy for each interface $\Omega_{\mathrm{if}}$
with interfacial bending rigidity $\kappa_{\mathrm{if}}$, and (total) mean curvature of the interface $H_{\mathrm{if}}$. In this model we penalize interfacial curvature energetically.
@@
Fig. 6(f) shows configurations of the defect cell in this BVM with bending rigidity for different apicobasal tensions $\Gamma$. For $\Gamma\leq 0.5$ we see that indeed the pentagonal cell is pushed out of the monolayer, only being attached because the lateral interfaces cannot be broken in the simulations. The luminal interfaces bend outward despite the energetic penalty to facilitate this extrusion process. For large $\Gamma$ we indeed also find a completely bulged state, as predicted by the continuum model.
@@
IV.3 Pentagonal and hexagonal defects in shells with curvature screening and lower defect tension
@@
For this configuration the total energy is now identical to the contributions from Eq. (24), but without the cone energy and with prescribed $R_{\mathrm{B}}$ and new angle $\gamma/2$, i.e.
@@
$E_{\mathrm{tot}}$ $=E_{\mathrm{core}}+E^{\mathrm{mean}}_{\mathrm{bending}}+E^{\mathrm{Gauss}}_{\mathrm{bending}}$ (28)
$\quad+E_{\mathrm{stretch}}+E_{\mathrm{bndry}}.$
Figure 7: Extrusion through lower defect tension in hexagons and pentagons. (a) Schematic depiction of defect in mean-field tissue with background curvature (radius $R_{\mathrm{bg}}$). (b) The energy difference for configurations with different defect tensions $\Gamma_{\mathrm{defect}}$ from the VM case with identical defect tension $\Gamma_{\mathrm{defect}}=\Gamma$. Solid line is local minimization from VM case, dashed line is extruded case. Red is pentagonal, black hexagonal defect cell. (c) Change of the critical defect tension $\Gamma_{\mathrm{defect}}^{\mathrm{crit}}$ at which the bulged state becomes unstable toward extrusion for different background radii $R_{\mathrm{bg}}$. (d) Difference of critical tensions in pentagonal and hexagonal cells in (c). Parameters: $\Gamma=0.9$; (b,c) $R_{\mathrm{bg}}/\tilde{h}=5$.
@@
We now also want to consider extrusion of a hexagonal cell, for which we have no azimuthal stretching, i.e. a disclinicity of $s=0$ in Eq. (19), and consider $n=6$ lateral faces. The geometric constants then read $\eta_{\mathrm{geom}}=3^{1/2}/2$ and $\zeta_{\mathrm{geom}}=3^{3/2}/2$. We have found that lower apico-basal tensions lead to extrusion, because the cell has to be completely engulfed by the apical interface in the extruded case. This suggests that one potential mechanism of extrusion is the lowering of apico-basal tension in the defect cell. Fig. 7(b) shows the energy differences from the VM reference for different defect tensions $\Gamma_{\mathrm{defect}}$ in a mean field tissue of tension $\Gamma=0.9$. We see that pentagonal energies are generally smaller in deformed states due to the relaxation of azimuthal strain when localizing Gaussian curvature in the defect.
@@
For both hexagons and pentagons we see that the extruded and bulged states become energetically more favorable as $\Gamma_{\mathrm{defect}}$ decreases. Below a critical tension $\Gamma_{\mathrm{defect}}^{\mathrm{crit}}$ the bulged state becomes unstable toward the extruded state. In Fig. 7(c) these critical tensions for pentagons and hexagons are plotted as functions of the background curvature radius $R_{\mathrm{bg}}/\tilde{h}$. For a small radius the critical tension is high, implying that extrusion is easy. The reason for this is that the volume constraint does not allow for a full relaxation of the Gaussian curvature in the defect, because we have assumed that $\delta/2<\pi/2-\gamma/2$, i.e. the cap region cannot accommodate negative curvature. This leads to a less stable bulged state and thus earlier instability toward extrusion. For intermediate background radii the bulged state is able to stabilize by relaxing the remaining Gaussian curvature in the defect. Increasing the background radius further then leads to less stability of the bulged state, because the possible opening angle is small anyways.
@@
In the pentagonal cell we see an intermediate state in Fig. 7(b): for intermediate tensions a local minimum of full bulging ($b=0$, $\delta/2>0$) forms, which then becomes unstable toward the extruded case for lower tensions (cf. Fig. 7(a)). In general we see that for decreasing defect tensions in both pentagons and hexagons the energy landscape loses local minima corresponding to bulged states, as the extruded case becomes increasingly more energetically favorable.
@@
Fig. 7(d) depicts the difference in the critical defect tensions of pentagons and hexagons, $\Gamma_{\mathrm{defect}}^{\mathrm{crit(5)}}-\Gamma_{\mathrm{defect}}^{\mathrm{crit(6)}}$. The necessary tension decrease for overcoming the energy barrier for extrusion in hexagons is approximately $10\%$ of the total tension higher than for pentagonal cells. While this suggests pentagonal cells to be more prone to extrusion, it also does not constitute a difference large enough for other cellular processes to not be able to overcome the barrier for hexagons as well.
IV.4 Effect of line tension on extrusion
Besides a reduction of apico-basal interfacial tension, another possible mechanism for extrusion are contractile actin structures pushing cells outward [3]. To consider such a process in our model, we add a line tension on the basal interface boundary, which effectively constitutes a contractile contribution along the basal interface edge. For a tension $\tau$ we thus have an energy
$E_{\mathrm{line}}=n\,\tau\,b.$ (29)
Effectively this constitutes an additional $b$-dependent energy, which leads to a tilting of the energy landscape in configuration space, because in the case of curvature screening no $R_{0}$ minimization is performed. Fig. 8 displays such energy landscapes for pentagons ($n=5$) for different line tensions $\tau$. We do see that the tilting can destabilize bulged states. However, since at $b=0$ the line tension does not contribute anymore, an energy barrier in the case of a transitioning from a bulged state with $b=0$ to the extruded case, is not changed by the line tension. In light of partially extruded cells, such as in Fig. 6(f), this implies that an additional tuning of interfacial tensions (either lateral or apico-basal) is necessary to achieve full extrusion.
IV.5 Instability modulation through luminal pressure and interface rigidity
Figure 9: Modulation of instability via luminal pressure and interfacial rigidity. (a) The shell from Fig. 1(c) is simulated with a luminal volume constraint $V/V_{\mathrm{VM}}=2$, corresponding to a luminal pressure. (b) Opening angles in the BVM for different volume constraints, where $V/V_{\mathrm{VM}}\approx 1$ corresponds to the case without a luminal volume constraint. (c) Rigidified shell with interfacial bending rigidity $\kappa_{\mathrm{if}}=0.7$. (d) Opening angles for different $\kappa_{\mathrm{if}}$, where $\kappa_{\mathrm{if}}=0$ is the BVM. Markers in (b,d) are average angles and shaded region shows $5\%$ to $95\%$ quantiles.
To this point the effects of pressure and bending stiffness have been neglected. Fig. 9(a) shows a shell whose luminal volume was prescribed, corresponding to an internal luminal pressure. Internal pressure is able to unfold pentagonal defect cells. Contrary to the different opening angles for different cellular symmetries, cf. Fig. 1(c), pressure leads to convergence of the average opening angles of pentagons, hexagons, and heptagons as it increases, see Fig. 9(a). In addition, we also see that the distribution of opening angles in pressurized BVM shells is less broad, more closely resembling the classical VM case. This means that pressure can be used to suppress the buckling instability caused by the surface tensions.
To highlight the role of interface curvature we added an interfacial bending energy to the BVM minimization, Eq. (25), with different interfacial bending rigidities $\kappa_{\mathrm{if}}$. An exemplary shell shown in Fig. 9(c) displays less curved interfaces and also the defect cells are unfolded. Fig. 9(d) shows the opening angles, whose values quickly approach each other as $\kappa_{\mathrm{if}}$ is increased. However, contrary to luminal pressure, rigidifying the interfaces does not achieve a full convergence of the opening angles toward each other.
This shows that luminal pressures and rigid interfaces decrease the importance of topological defects in the BVM, which could serve as a mechanism for cells to achieve symmetry breaking and control extrusion and cell geometry. Moreover, including interfacial curvatures, as is necessary to describe some experimental systems, changes the importance of such defects drastically. In theoretical models of tissues it is thus important to reconsider the role of defects and the exact formulation of models, potentially rendering curvature energies, which are often neglected, central to capture single-cell behavior.
V Discussion

### v2

@@ III.3 Extrusion of a bubbly defect cell in a mean-field vertex model tissue at icosahedral tips
In the following we assume the mean-field tissue parameters to be from the classical (non-bubbly) VM, thus only allowing for bulging of the defect cell. In principle, however, mean-field parameters could be derived for the BVM in a similar fashion as for the VM, or could be taken from experimental measurements for the tissue properties.
@@
Figure 8: Energy landscapes in configuration space and the resulting defect cell shapes. (a-c) Energy landscapes with coordinates $(\delta/2,b)$ for different surface tensions $\Gamma$. The black line marks the vertex model limit without bulging of the defect cell and the red circle marks the minimum in this limit. The dashed line is the projection onto the $(\Delta E,\delta/2)$ plane of this limit. The color codes the energy difference $\Delta E$ with respect to the reference energy, which is the vertex model minimum. The corners of the landscape at $(\delta/2,b)=(0,0)$ and $(\delta_{\mathrm{max}}/2,0)$ correspond to the extruded and maximally bent situations, respectively, cf. Fig. 6. (d) Continuum model energy differences from the VM reference $\Delta E$ of the minimal energy configuration (solid), starting from the VM reference, and of the extruded configuration (dashed). (e,f) Corresponding defect opening angles $\delta/2$ and outer interfacial area $A_{\mathrm{outer}}$. (g) For the BVM with interfacial bending rigidity the defect cell is extruded for small apico-basal tension $\Gamma=0.5$, and partially extruded for large tensions $\Gamma=0.7$ and $\Gamma=1.0$. (h,i) Corresponding opening angles and outer interfacial area for the BVM. Shells of Caspar-Klug size $(T_{1},T_{2})=(12,0)$ with bending rigidity $\kappa_{\mathrm{if}}=5$ in dimensionless units.
@@
The total energy is the sum of the contributions from the defect, the cap and the outer cone region. We consider the energy difference to a reference state to quantify the energy release due to bulging by considering the energy difference $\Delta E$ from the vertex model case with minimal energy. The latter corresponds to the full defect model described in Appendix G, see Fig. 12.
@@
We find that for all configurations characterized by $\delta/2$ and $a$ the energy has a unique minimum in the tissue buckling radius $R_{\mathrm{B}}$. We assume this radius to be minimal for every defect cell shape and thus can describe all possible shapes by the variables $(\delta/2,a)$, which constitutes the configuration space. Both parameters are bounded from below by 0, where $(0,0)$ is possible and corresponds to the fully extruded case, and from above by the volume constraint of the defect cell.
@@
Fig. 8(a,b,c) show energy landscapes in configuration space for different surface tensions $\Gamma$. For now we assume identical cell properties, i.e. the tension in the defect cell matches the one of the surrounding tissue. We notice that, by bulging outward at the defect, the energy of the tissue can be reduced considerably in comparison to the classical vertex model reference case. For small $\Gamma$ more energy can be released by bulging outward than for larger $\Gamma$. For $\Gamma>0.5$ we observe that two minima form in the energy, one in the extruded case $(\delta/2,a)=(0,0)$, and one in a bulged case with very small $a$. The energy minimum in extrusion is higher than the bulged one for larger $\Gamma$ and an energy barrier has to be overcome from both the reference and the global minimum configuration to extrude the defect.
@@
We now numerically determined the (local) minimal energy configurations for different $\Gamma$, starting from the VM reference state. Fig. 8(d,e,f) depict the resulting minimal energy configurations. Above a critical tension $\Gamma$ the luminal (basal) interface collapses but the angle remains finite, indicating a completely bulged state. The fully extruded state at $(\delta/2,a)=(0,0)$ always represents a possible configuration, but its energy increases with $\Gamma$. Looking at the energy difference from the VM state, a configuration with interfacial curvature is always energetically more favorable, cf. Fig. 8(e,f). This is not surprising as the surface area can be minimized by the cell becoming more spherical. The extruded case, on the other hand, requires more energy for large apico-basal tension, as we need to wrap the cell in the energetically (relatively) expensive apical interface. Lowering the tension allows for an energy release in the extruded case, which can become lower than in the bulged state. Note, however, that a bulged state persists beyond the point at which the extruded case is energetically more favorable and that an energy barrier exists in configuration space, which prevents the cell from being directly extruded from the reference VM configuration, cf. Fig. 8(b).
@@
Considering icosahedral shells in the BVM and adding a bending rigidity across triangulation points of hexagonal cells’ interfaces enables a numerical simulation similar to our continuum approach of a bubbly defect cell surrounded by VM cells, cf. Fig. 4(g). We thus consider an additional bending energy, Eq. (3).
@@
Fig. 8(g) shows configurations of the defect cell in this BVM with bending rigidity for different apicobasal tensions $\Gamma$. For $\Gamma\leq 0.5$ we see that indeed the pentagonal cell is pushed out of the monolayer, only being attached because the lateral interfaces cannot be broken in the simulations. This is also visible in the the parameters quantifying the configurations in Fig. 8(h,i) The luminal interfaces bend outward despite the energetic penalty to facilitate this extrusion process.
For intermediate $\Gamma$ the neighbor cells are deformed due to finite $\kappa_{\mathrm{if}}$, while for large $\Gamma$ a completely bulged state is found, as predicted by the continuum model.
@@
III.4 Extrusion at pentagonal and hexagonal defects in spherical shells with curvature screening
Figure 9: Extrusion through lower defect tension in hexagons and pentagons. (a) Schematic depiction of defect in mean-field tissue with background curvature (radius $R_{\mathrm{bg}}$). (b) The energy difference for continuum model configurations with different defect tensions $\Gamma_{\mathrm{defect}}$ from the VM case with identical defect tension $\Gamma_{\mathrm{defect}}=\Gamma$. Solid line is local minimization from VM case, dashed line is extruded case. Red is pentagonal, black hexagonal defect cell. (c) BVM-shells from Fig. 1 in which pentagonal cells have different apico-basal tensions $\Gamma_{\mathrm{defect}}^{(5)}$, displaying partially extruded morphologies that depend on the tension. Pentagons in red, heptagons in blue. (d) Cross section of a pentagonal cell surrounded by hexagonal cells being extruded upon lowering of $\Gamma_{\mathrm{defect}}^{(5)}$ in the BVM. (e) Average outer interfacial area of BVM-shells; horizontal dashed line corresponds to fully extruded cells. (f) Change of the critical defect tension $\Gamma_{\mathrm{defect}}^{\mathrm{crit}}$ in the continuum model. at which the bulged state becomes unstable toward extrusion for different background radii $R_{\mathrm{bg}}$. (g) Difference of critical tensions in pentagonal and hexagonal cells in (f). Parameters: $\Gamma=0.9$; (b,f) $R_{\mathrm{bg}}/\tilde{h}=5$.
@@ III.4 Extrusion at pentagonal and hexagonal defects in spherical shells with curvature screening
For this configuration the total energy is now identical to the contributions described in Appendix G (cf. Eq. (G)), but without the cone energy and with prescribed $R_{\mathrm{B}}$ and new angle $\gamma/2$, i.e.
@@
$aligned E_{\mathrm{tot}}&=E_{\mathrm{core}}+E^{\mathrm{mean}}_{\mathrm{bending}}+E^{\mathrm{Gauss}}_{\mathrm{bending}}\\ &\quad+E_{\mathrm{stretch}}+E_{\mathrm{bndry}}.$ (10)
@@
We now also want to consider extrusion of a hexagonal cell, for which we have no azimuthal stretching (a disclinicity of $s=0$ in Eq. (29) in Appendix F) and consider $n=6$ lateral faces. The geometric constants then read $\eta_{\mathrm{geom}}=3^{1/2}/2$ and $\zeta_{\mathrm{geom}}=3^{3/2}/2$. We have found that lower apico-basal tensions lead to extrusion, because the cell has to be completely engulfed by the apical interface in the extruded case. This suggests that one potential mechanism of extrusion is the lowering of apico-basal tension in the defect cell. Fig. 9(b) shows the energy differences from the VM reference for different defect tensions $\Gamma_{\mathrm{defect}}$ in a mean field tissue of tension $\Gamma=0.9$. We see that pentagonal energies are generally smaller in deformed states due to the relaxation of azimuthal strain when localizing Gaussian curvature in the defect.
For both hexagons and pentagons we see that the extruded and bulged states become energetically more favorable as $\Gamma_{\mathrm{defect}}$ decreases. Below a critical tension $\Gamma_{\mathrm{defect}}^{\mathrm{crit}}$ the bulged state becomes unstable toward the extruded state.
@@
In BVM-simulations, we find strong outward bulging reminiscent of intermediate states in cell extrusion upon lowering of the apical and basal defect tension, without achieving full extrusion due to the inability of closing the forming hole in the simulations (Fig. 9(c,d,e)). In Fig. 9(f) these critical tensions for pentagons and hexagons are plotted as functions of the background curvature radius $R_{\mathrm{bg}}/\tilde{h}$. For a small radius the critical tension is high, implying that extrusion is easy. The reason for this is that the volume constraint does not allow for a full relaxation of the Gaussian curvature in the defect, because we have assumed that $\delta/2<\pi/2-\gamma/2$, i.e. the cap region cannot accommodate negative curvature. This leads to a less stable bulged state and thus earlier instability toward extrusion. For intermediate background radii the bulged state is able to stabilize by relaxing the remaining Gaussian curvature in the defect. Increasing the background radius further then leads to less stability of the bulged state, because the possible opening angle is small anyways.
@@
In the pentagonal cell we see an intermediate state in Fig. 9(b): for intermediate tensions a local minimum of full bulging ($a=0$, $\delta/2>0$) forms, which then becomes unstable toward the extruded case for lower tensions (cf. Fig. 8(a,b)). In general we see that for decreasing defect tensions in both pentagons and hexagons the energy landscape loses local minima corresponding to bulged states, as the extruded case becomes increasingly more energetically favorable.
@@
Fig. 9(g) depicts the difference in the critical defect tensions of pentagons and hexagons, $\Gamma_{\mathrm{defect}}^{\mathrm{crit(5)}}-\Gamma_{\mathrm{defect}}^{\mathrm{crit(6)}}$. The necessary tension decrease for overcoming the energy barrier for extrusion in hexagons is approximately $10\%$ of the total tension higher than for pentagonal cells. While this suggests pentagonal cells to be more prone to extrusion, it also does not constitute a difference large enough for other cellular processes to not be able to overcome the barrier for hexagons as well. Introducing additional luminal line tensions yields a destabilization of a bulged state with a finite size luminal interface, does not contribute, however, to the extrusion process without an involvement of lateral tensions, see Appendix E.
IV Discussion